OpenStudy (anonymous):

could someone please show me how to find the product for: (x+3)^2

OpenStudy (anonymous):

Distribute the terms, since what you really have is \((x+3)(x+3)\).\[\Large (\color{red} a + \color{red} b)(\color{blue}c+\color{blue}d)=\color{red}a\color{blue}c+\color{red}a\color{blue}d+\color{red}b\color{blue}c+\color{red}b\color{blue}d\]

OpenStudy (anonymous):

No...it's going to be a quadratic expression.

OpenStudy (anonymous):

\[\Large (x+3)^2=(\color{red} x + \color{red} 3)(\color{blue}x+\color{blue}3)=\cdots\]

OpenStudy (anonymous):

No. Follow the pattern. \[\Large (x+3)^2=(\color{red} x + \color{red} 3)(\color{blue}x+\color{blue}3)=\color{red}x\color{blue}x+\color{red}x\cdot \color{blue}3+\color{red}3\cdot \color{blue}x+\color{red}3\cdot \color{blue}3\]

OpenStudy (anonymous):

x^4+12

OpenStudy (anonymous):

No. Simplify what I wrote.